Two cars start from the same place traveling in opposite directions. One car travels 4 miles per hours faster than the other car. Find the speed their speeds if after 5 hours they are 520 miles apart.
Answer by checkley77(12844)

5x+5(x+4)=520
5x+5x+20=520
10x=520-20
10x=500
x=500/10
x=50 mph for the slower car.
50+4=54 mph for the faster car.
Proof:
5*50+5*54=520
250+270=520
5290=520